OpenStudy (anonymous):

Please help, i tried twice already. What is the acceleration of an object if it goes from a velocity of +25 m/s to rest in 5.0 s? A. –5 m/s2 B. +5 m/s2 C. –25 m/s2 D. +25 m/s2

OpenStudy (anonymous):

\[v(t)=v_o+at\] \[v(5)=0\] \[0 = 25 + a \times 5\] \[5a = -25\] \[a = -5_{m/s^2}\]

jimthompson5910 (jim_thompson5910):

Go here http://www.studyphysics.ca/newnotes/20/unit01_kinematicsdynamics/chp04_acceleration/lesson11.htm scroll down to where it says `"How to Choose the Right Formula!"` (almost near the bottom) you'll find this table http://www.studyphysics.ca/newnotes/20/unit01_kinematicsdynamics/chp04_acceleration/images/formulas.GIF it says `"To choose the correct formula, look for the one thing that you are not given or asked for as the 'missing' part of a formula."` We're not given the distance and we don't want to find the distance. So we look for an X under the d column and we find that we have to use this formula \[\Large V_{f} = V_{i} + a*t\]

jimthompson5910 (jim_thompson5910):

In this case, \[\Large V_{f} = 0, \ V_{i} = 25, \ t = 5\]